# Reflectionless and echo modes in asymmetric Damour-Solodukhin wormholes

### v1

@@ Abstract
It is understood that the echo waveforms in ultracompact objects can be regarded as composed mainly of the asymptotic high-overtone quasinormal modes, dubbed echo modes, which predominantly lie parallel to the real frequency axis. Alternatively, Rosato et al. recently suggested that high-frequency quasi-reflectionless scattering modes are primarily responsible for the echo phenomenon. This identification relies on greybody factors as stable observables, despite the apparent spectral instability of quasinormal modes. In this work, by extending the definition of quasi-reflectionless modes to reflectionless ones and generalizing symmetric Damour-Solodukhin wormholes to asymmetric cases, we examine the underlying similarity between the reflectionless and echo mode spectra in the complex frequency plane. Through a primarily analytical treatment, we demonstrate that the asymptotic properties of these two spectra exhibit a strong resemblance, featuring an approximately uniform distribution parallel to the real frequency axis with the same spacing between successive modes. Specifically, the real parts of echo modes coincide with those of reflectionless modes at the limit $|\mathrm{Re}\omega|\gg|\mathrm{Im}\omega|$. While echo modes typically possess non-vanishing imaginary parts, the reflectionless modes of symmetric Damour-Solodukhin wormholes lie precisely on the real frequency axis, with any deviation serving as a measure of the degree of asymmetry of the wormhole. We support our derivations by employing two complementary approaches, based on the scattering matrix and the Green’s function, and argue that both perspectives provide effective tools for describing the echo phenomenon.
@@ I Introduction
The significance of spectral instability resides in its observational implications, especially regarding black hole spectroscopy [19]. The relevance of the topic is closely connected with the detections of gravitational waves emanating from the binary mergers by the LIGO and Virgo collaboration [20, 21, 22, 23]. The success of the ground-based facilities has further promoted the ongoing spaceborne projects, such as LISA [24], TianQin [25, 26], and Taiji [27], leading to speculation that direct observation of ringdown waveforms might also be plausible [28]. Under realistic astrophysical conditions, the source of gravitational waves is not a single isolated celestial object. Typically, they are submerged and interacting with surrounding matter. This leads to deviations from the ideal metric with perfect symmetries, motivating the studies of the so-called “dirty” black holes [29, 30, 31, 32] in the context of black hole perturbation theory. The obtained results on spectral stability warrant further scrutiny of these dirty compact objects. Regarding the deformed spectrum in the frequency domain, the asymptotic QNMs that align almost parallel to the real axis are closely related to the intriguing concept of echoes, a late-time phenomenon first speculated by Cardoso et al. [15, 33]. As potential observables, echoes were proposed to help distinguish different but otherwise similar gravitational systems via their distinct properties near the horizon. Such a physical picture has motivated many studies into echoes across various systems, including exotic compact objects such as gravastars [34, 35], boson stars [36], and wormholes [37, 38, 39, 40]. Similar to the late-time tail, echoes are also attributed to the analytic properties of Green’s function, as analyzed by Mark et al. [41]. In studies of Damour-Solodukhin type wormholes [42], Bueno et al. [43] have investigated echoes by explicitly solving for specific frequencies at which the transfer matrix becomes singular, providing further insights into the complex interplay of spacetime geometry and QNMs. It has been argued [44] that the spectral instability is closely associated with the black hole echoes, based on the frequency-domain analyses [16, 45, 46].
@@
The related topic of the spectral instability, echoes, and causality has incited many studies [47, 18, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62]. Notably, Cheung et al. [49] pointed out that even the fundamental mode can be destabilized under rather generic perturbations. By introducing a small perturbation to the Regge-Wheeler effective potential, it was shown [49] that the fundamental mode undergoes an outward spiral while the deviation’s magnitude increases. In Refs. [63, 64], the numerical findings on the instability of the fundamental mode were interpreted in terms of the instability of an approximated case where the metric perturbation is modeled by an insignificant disjoint potential barrier. However, if one does not assume that the perturbation is disjoint, some subtlety emerges. In particular, it was demonstrated [65] that, contrary to the existing results, the fundamental mode might be stable, whose validity is confirmed by numerical calculations without introducing any approximation. This is in agreement with previous findings [65], reinforcing that the low-lying modes are indeed more resilient than the high overtones.
@@
More recently, the notion of quasi-reflectionless scattering modes (RSMs) was introduced [76]. The proposal is motivated by the connection between the greybody factors and the ringdown signal. In particular, the authors explored symmetric cavity potentials, which mimic the case of a wormhole. It was shown that these high-frequency quasi-RSMs are primarily responsible for the echoes of the underlying compact object. These modes were found to be uniformly distributed in the frequency domain, parallel to the real frequency axis.
Intriguingly, the above properties are reminiscent of those of echo modes discussed in [16, 46, 44], given that their definitions are rather distinct. The present study is thus motivated and aims to explore the underlying similarity between the quasi-reflectionless and echo modes. By promoting the definition of the quasi-RSMs to (entirely) reflectionless ones, these frequencies are defined by complex numbers, corresponding to the poles of appropriately modified Green’s function. Besides, we explore asymmetric cavity potentials, specifically asymmetric Damour-Solodukhin wormholes, and demonstrate that the reminiscence in the asymptotic properties of the reflectionless and echo mode spectra can be largely derived analytically. In particular, we show that the RSMs in asymmetric Damour-Solodukhin wormholes naturally inherit the quasinormal modes of the metric situated on one side of the throat, as well as the RSMs from the other side. We carry out the derivations using two different approaches, based on the evaluation of the scattering amplitudes and the poles of the modified Green’s function.
The remainder of this paper is organized as follows. In Sec. II, we briefly review Green’s function approach to the black hole QNMs and show that it is readily generalized to the case of reflectionless scattering processes. Subsequently, in Sec. III, we derive the RSMs as complex frequencies using the scattering amplitude for asymmetric Damour-Solodukhin wormholes. The properties and implications of the solutions are discussed and compared against the echo modes. An alternative derivation for RSMs is presented in Sec. IV, using the modified Green’s function. The reminiscence in the two different scenarios is demonstrated in a more transparent fashion. In Sec. V, we illustrate our results by elaborating on a few semi-analytically assessable examples. The concluding remarks are given in the last section. The complementary mathematical derivations will be relegated to Appx. AB, and C.
@@ III Reflectionless scattering modes and their asymptotic properties
As mentioned, the RSMs can be defined as the analytic continuation of quasi-reflective modes, achieved by promoting the frequency from the real axis to the complex plane. They correspond to the complex frequencies that asymptotically satisfy the boundary condition Eq. (13). Similar to the QNMs, this definition can be implemented either in terms of the transfer matrix $\mathbb{T}$, when viewed as a scattering problem, or in terms of the poles of the Green’s function, when one considers the dynamical evolution of initial perturbations.
@@ IV An altnernative approach based on Green’s function
V A few numerical examples
In this section, we elaborate on a few examples to illustrate the results obtained in the previous sections.
The first example is the double-$\delta$ potential of the form
@@ V A few numerical examples
Symmetric case 4.33175 7.56372 10.7587 180.612 183.754 186.896
-2.23338ii -2.75924ii -3.09718ii -5.88994ii -5.90718 $i$ -5.92412ii
Est. by Eq. (68) 4.71239 7.85398 10.9956 180.642 183.783 186.925
-2.24334ii -2.75417ii -3.09064ii -5.88966ii -5.9069ii -5.92385ii
Asymmetric case 4.46629 7.64109 10.8133 180.615 183.757 186.899
-1.8752ii -2.40346ii -2.74432ii -5.54329ii -5.56053ii -5.57748ii
Est. by Eq. (68) 4.71239 7.85398 10.9956 180.642 183.783 186.925
-1.89677ii -2.40759ii -2.74407ii -5.54309ii -5.56033ii -5.57728ii
@@ VI Concluding remarks
In conclusion, this work explored the concept of RSMs in the context of asymmetric Damour-Solodukhin wormholes. By promoting from the real axis to the complex plane, we elaborate on the underlying similarity between the reflectionless and echo mode spectra. For asymptotic modes satisfying $|\mathrm{Re}\omega|\gg|\mathrm{Im}\omega|$, it is shown that these two spectra exhibit a strong resemblance, featuring an approximately uniform distribution parallel to the real frequency axis with the same spacing between successive modes. In particular, it was shown that the real parts of asymptotic QNMs coincide with those of RSMs. While echo modes typically possess non-vanishing imaginary parts, the RSMs of Damour-Solodukhin wormholes mainly lie close to the real frequency axis, where the distance between the RSMs and the real frequency axis measures the degree of deviation from a perfectly symmetric Damour-Solodukhin. Compared with echo modes, this property of RSMs leads to a more substantial impact on the amplitudes of the time-domain waveforms, and subsequently the echoes. We carried out the derivations by employing two complementary approaches, based on the scattering matrix and the Green’s function.
@@ Acknowledgements
Appendix A Flux conservation and the sign of the imaginary parts of the reflectionless modes

### v2

@@ Abstract
It is understood that the echo waveforms in ultracompact objects can be regarded as composed mainly of the asymptotic high-overtone quasinormal modes, dubbed echo modes, which predominantly lie parallel to the real frequency axis. Alternatively, Rosato et al. recently suggested that high-frequency quasi-reflectionless scattering modes are primarily responsible for the echo phenomenon. This identification relies on greybody factors as stable observables, despite the apparent spectral instability of quasinormal modes. In this work, by extending the definition of quasi-reflectionless modes to reflectionless ones and generalizing symmetric Damour-Solodukhin wormholes to asymmetric cases, we examine the underlying similarity between the reflectionless and echo mode spectra in the complex frequency plane. Through a primarily analytical treatment, we demonstrate that the asymptotic properties of these two spectra exhibit a strong resemblance, featuring an approximately uniform distribution parallel to the real frequency axis with the same spacing between successive modes. Specifically, the real parts of echo modes coincide with those of reflectionless modes at the limit $|\mathrm{Re}\omega|\gg|\mathrm{Im}\omega|$. While echo modes typically possess non-vanishing imaginary parts, the reflectionless modes of symmetric Damour-Solodukhin wormholes lie precisely on the real frequency axis, with any deviation serving as a measure of the degree of asymmetry of the wormhole. We support our derivations by employing two complementary approaches, based on the scattering matrix and the Green’s function. For a given identical source, the waveforms are calculated numerically using the Green’s functions. The amplitudes of the waveforms associated with reflectionless modes are found to be more pronounced than those of the echo modes, because reflectionless modes typically lie closer to the real frequency axis than the latter. It is argued that both perspectives provide effective tools for describing the echo phenomenon.
@@ I Introduction
The significance of spectral instability resides in its observational implications, especially regarding black hole spectroscopy [19]. The relevance of the topic is closely connected with the detections of gravitational waves emanating from the binary mergers by the LIGO and Virgo collaboration [20, 21, 22, 23]. The success of the ground-based facilities has further promoted the ongoing spaceborne projects, such as LISA [24], TianQin [25, 26], and Taiji [27], leading to speculation that direct observation of ringdown waveforms might also be plausible [28]. Under realistic astrophysical conditions, the source of gravitational waves is not a single isolated celestial object. Typically, they are submerged and interacting with surrounding matter. This leads to deviations from the ideal metric with perfect symmetries, motivating the studies of the so-called “dirty” black holes [29, 30, 31, 32] in the context of black hole perturbation theory. The obtained results on spectral stability warrant further scrutiny of these dirty compact objects. Regarding the deformed spectrum in the frequency domain, the asymptotic QNMs that align almost parallel to the real axis are closely related to the intriguing concept of echoes, a late-time phenomenon first speculated by Cardoso et al. [15, 33]. As potential observables, echoes were proposed to help distinguish different but otherwise similar gravitational systems via their distinct properties near the horizon. Such a physical picture has motivated many studies into echoes across various systems, including exotic compact objects such as gravastars [34, 35], boson stars [36], and particularly wormholes [37, 38, 39, 40, 41, 42, 43, 44, 45]. Similar to the late-time tail, echoes are also attributed to the analytic properties of Green’s function, as analyzed by Mark et al. [46]. In studies of Damour-Solodukhin type wormholes [47], Bueno et al. [48] have investigated echoes by explicitly solving for specific frequencies at which the transfer matrix becomes singular, providing further insights into the complex interplay of spacetime geometry and QNMs. It has been argued [49] that the spectral instability is closely associated with the black hole echoes, based on the frequency-domain analyses [16, 50, 51].
@@
The related topic of the spectral instability, echoes, and causality has incited many studies [52, 18, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69]. Notably, Cheung et al. [54] pointed out that even the fundamental mode can be destabilized under rather generic perturbations. By introducing a small perturbation to the Regge-Wheeler effective potential, it was shown [54] that the fundamental mode undergoes an outward spiral while the deviation’s magnitude increases. In Refs. [70, 71], the numerical findings on the instability of the fundamental mode were interpreted in terms of the instability of an approximated case where the metric perturbation is modeled by an insignificant disjoint potential barrier. However, if one does not assume that the perturbation is disjoint, some subtlety emerges. In particular, it was demonstrated [72] that, contrary to the existing results, the fundamental mode might be stable, whose validity is confirmed by numerical calculations without introducing any approximation. This is in agreement with previous findings [72], reinforcing that the low-lying modes are indeed more resilient than the high overtones.
@@
More recently, the notion of quasi-reflectionless scattering modes (quasi-RSMs) was introduced in Ref. [83]. This proposal is motivated by the connection between greybody factors and ringdown waveforms. For symmetric cavity potentials, which mimic wormhole-like spacetime configurations, perfect RSMs exist at discrete real frequencies, whereas in more general situations the minima of the reflectivity correspond to quasi-RSMs. As an intriguing feature of the greybody factor in the high-frequency regime, these quasi-RSMs were shown to be primarily responsible for the echo phenomenon of the underlying compact object. These modes were found to be approximately uniformly distributed along the frequency axis.
Intriguingly, the above properties are reminiscent of those of echo modes discussed in [16, 51, 49], even though their definitions are rather different. The present study is therefore motivated to explore the underlying similarity between quasi-RSMs and echo modes. By promoting the definition of quasi-RSMs to fully reflectionless modes, these frequencies become complex and correspond to the poles of an appropriately modified Green’s function. In addition, we investigate asymmetric cavity potentials, specifically asymmetric Damour-Solodukhin wormholes, and demonstrate that the resemblance in the asymptotic properties of the reflectionless and echo-mode spectra can be largely obtained analytically. We carry out explicit derivations of the spectra of RSMs and echo modes using two different approaches, based on the evaluation of the scattering amplitudes and on the poles of the modified Green’s function. In general, the RSMs are found to lie closer to the real frequency axis, leading to more pronounced waveforms associated with RSMs than those related to echo modes for a given identical source.
The remainder of this paper is organized as follows. In Sec. II, we briefly review Green’s function approach to the black hole QNMs and show that it is readily generalized to the case of reflectionless scattering processes. Subsequently, in Sec. III, we derive the RSMs as complex frequencies using the scattering amplitude for asymmetric Damour-Solodukhin wormholes. The properties and implications of the solutions are discussed and compared against the echo modes. An alternative derivation for RSMs is presented in Sec. IV, using the modified Green’s function. The reminiscence in the two different scenarios is demonstrated in a more transparent fashion. In Sec. V, we illustrate our results by elaborating on a few semi-analytically tractable examples. In particular, the waveforms associated with RSMs and echo modes are evaluated and analyzed. The concluding remarks are given in the last section. The complementary mathematical derivations and tecnical details will be relegated to Appx. ABCD and E.
@@ III Reflectionless scattering modes and their asymptotic properties
As mentioned, the RSMs can be defined as the analytic continuation of quasi-reflective modes, achieved by promoting the frequency from the real axis to the complex plane. They correspond to the complex frequencies that asymptotically satisfy the boundary condition Eq. (13). Similar to the QNMs, this definition can be implemented either in terms of the transfer matrix $\mathbb{T}$, when viewed as a scattering problem, or in terms of the poles of the Green’s function, when one considers the dynamical evolution of some initial conditions.
@@ IV An altnernative approach based on Green’s function
V Numerical examples
In this section, we present a few examples to illustrate the results obtained in the previous sections. First, in Sec. V.1, we evaluate and compare the spectra of quasi-RSMs, RSMs, and echo modes in two toy-model wormholes. Subsequently, in Sec. V.2 we compute and compare the corresponding time-domain waveforms for a given source term, exploring the differences that are closely connected to the underlying pole structures.
V.1 The spectra of quasi-RSMs, RSMs, and echo modes
We consider two toy-model wormholes. The first example is the double-$\delta$ potential of the form
@@ V.1 The spectra of quasi-RSMs, RSMs, and echo modes
Symmetric case 1.21427 4.33175 7.56372 180.612 183.754 186.896
-0.95224ii -2.23338ii -2.75924ii -5.88994ii -5.90718 $i$ -5.92412ii
Est. by Eq. (68) 1.57080 4.71239 7.85398 180.642 183.783 186.925
-1.14473ii -2.24334ii -2.75417ii -5.88966ii -5.9069ii -5.92385ii
Asymmetric case 1.54318 4.46629 7.64109 180.615 183.757 186.899
-0.79374ii -1.8752ii -2.40346ii -5.54329ii -5.56053ii -5.57748ii
Est. by Eq. (68) 1.57080 4.71239 7.85398 180.642 183.783 186.925
-0.79816ii -1.89677ii -2.40759ii -5.54309ii -5.56033ii -5.57728ii
@@
Before closing this subsection, it is instructive to directly compare the RSM spectrum with that of the QNMs. From the examples elaborated above, it is observed that the asymptotic modes in these cases have identical spacing along the real frequency axis, as analytically assessed in the previous sections. However, the RSM spectrum lies closer to the real axis than the QNMs, as demonstrated in Tabs. 1 and 2. While this feature guarantees identical echo periods in both scenarios, it leads to a nontrivial difference in the resulting waveforms. As will be elaborated in the following subsection, for a given identical source, the waveforms associated with RSMs are more pronounced than their counterparts related to QNMs.
@@
V.2 The waveforms of RSMs and echo modes
We now elaborate on the similarities and differences between the RSMs and echo modes through their time-domain waveforms. In this subsection, the waveforms are computed using Green’s functions constructed using Eqs. (7) and (14). The primary difference between them arises from the distinct boundary conditions imposed on $f(\omega,x)$ and $\widetilde{f}(\omega,x)$.
We consider asymmetric Damour-Solodukhin wormholes composed of two delta-function effective potential barriers with unequal magnitudes, adopting the same parameters as in the preceding subsection for the spectral calculations.
Using the explicit forms Eqs. (101) and (102) derived in Appx. E, the numerical results for the Green’s functions are shown in Fig. 7. It is apparent that the two Green’s functions differ significantly from each other. Meanwhile, unlike those observed with the Regge-Wheeler potential (cf. Fig. 1 of [50]), they both exhibit strong oscillations in the spatial and frequency domains. This behavior stems from the poles lying close to the real frequency axis, as demonstrated in the preceding subsection.
Refer to caption Refer to caption
Refer to caption Refer to caption
Fig. 7: The frequency-domain Green’s functions, Eqs. (7) and (14), for echo modes and RSMs of asymmetric Damour-Solodukhin wormholes composed of two delta-function effective potential barriers with different magnitudes. The calculations are performed using the same parameters as in Fig. 3, with $V_{0}^{\mathrm{L}}=2$, $V_{0}^{\mathrm{R}}=1$, and $x_{c}=\frac{1}{2}$. Upper row: The real part of the Green’s functions as a function of $\omega$ and $x$, with $x^{\prime}=y_{0}=-\frac{1}{4}$. Bottom row: The real and imaginary parts of the Green’s functions on the spatial slice $x=\frac{1}{4}$. The results associated with the RSMs are shown in the left column, while those corresponding to the echo modes are displayed in the right column.
We now compute the waveforms. To enable a meaningful comparison between the waveforms for RSMs and echo modes, for both cases, we adopt the following identical source in the frequency domain:
$S(\omega,y)=\sigma_{0}\delta(y-y_{0})\frac{\omega^{2}}{(\omega-\omega_{0})^{3}},$ (78)
where $\sigma_{0}=1$, $y_{0}=-\frac{1}{4}$, and $\omega_{0}=-4+2i$. This represents a pulse localized at $y=y_{0}$, with a spread in the frequency domain. Physically, this source term encodes the information about the initial/boundary conditions from the Fourier [90] and Laplace [86, 8] transform perspectives. As detailed further in Appx. E, the numerator $\omega^{2}$ suppresses the pole at the origin, the denominator ensures faster convergence of the inverse Fourier transform, and the additional pole at $\omega=\omega_{0}$ plays no role since the contour integration follows a large circle in the lower half of the complex frequency plane. The resulting frequency-domain waveform is given by
$\Psi(\omega,x)=\int_{y}G(\omega,x,y)S(\omega,y)dy=\sigma_{0}G(\omega,x,y_{0})\frac{\omega^{2}}{(\omega-\omega_{0})^{3}},$ (79)
which can be transformed into the time domain via the inverse Fourier transform.
The resulting time-domain waveforms are presented in Fig. 8. Unlike the Green’s functions, the time-domain waveforms closely resemble one another. As expected, both waveforms exhibit echoes with period $T=4x_{c}$, corresponding to the round-trip time between the two delta peaks. This is also consistent with $T={2\pi}/{\Delta\omega}$, where $\Delta\omega$} is the spacing along the real frequency axis between successive asymptotic modes governed by Eqs. (64) and (68), namely, $\Delta\omega={\pi}/{2x_{c}}$. As explained in Appx. E, one can verify the correctness of these waveforms by extracting the underlying modes using the Prony method [91]. For the echo waveforms, the dominant mode $\omega_{0}=1.54266-0.793539i$ is found, in reasonable agreement with the value 1.543180.79374i1.54318-0.79374i listed in Tab. 1. For the RSM waveforms, one instead obtains $\omega_{0}=1.92975-0.261073i$, consistent with the value 1.922850.260979i1.92285-0.260979i given in Tab. 1. In the present setup, owing to the toy-model nature of the potential, the echoes do not appear as an enveloping modulation with a shorter oscillation period primarily governed by the fundamental mode. This is because the oscillation period of the above extracted fundamental modes is much larger than the echo period, and thus is not visible in Fig. 8. It is worth noting that the echoes do not correspond to any individual mode. This is also demonstrated by the Prony fit, since no mode is found with real part $\mathrm{Re}\,\omega\approx 2\pi/T\approx 6.28$. As emphasized in Refs. [59, 16, 51, 92], the echoes arise as a collective effect due to a sequence of modes evenly distributed along the real frequency axis.
When comparing the waveforms associated with the RSMs to those of the QNMs, one observes that the reflectionless scattering waves exhibit larger amplitudes than the QNM contributions. In the present case, the magnitude of the RSM waveforms exceeds that of the echo modes by about one order of magnitude. This behavior is understood to be general, since the RSM spectrum lies closer to the real frequency axis than that of the echo modes, as clearly demonstrated in Tab. 1. As a result, despite their similarity, RSMs have a more significant impact on the time-domain waveforms.
Refer to caption Refer to caption
Fig. 8: The time-domain waveforms for asymmetric Damour-Solodukhin wormholes composed of two delta-function effective potential barriers with unequal magnitudes. The waveforms at $x=x_{0}=\frac{1}{4}$ are computed using the Green’s functions for RSM and echo modes shown in Fig. 7, for a source term given by Eq. (78). The results for the RSMs appear in the left panel, while those for the echo modes are shown in the right panel. The real and imaginary parts of the waveforms $\Psi$ are plotted versus time on a logarithmic scale, clearly revealing the “echo” feature, while the insets show the waveforms on the original linear scale. The echo period, visible in all four panels, corresponds to the spacing between successive asymptotic modes along the real frequency axis: $T=\frac{2\pi}{\Delta\omega}=4x_{c}$. The reflectionless scattering waves exhibit greater amplitude than those associated with the QNMs. As discussed in the text, this arises because the RSMs lie closer to the real frequency axis than the QNMs.
@@ VI Concluding remarks
In conclusion, this work explored the concept of RSMs in the context of asymmetric Damour-Solodukhin wormholes. By promoting from the real axis to the complex plane, we elaborate on the underlying similarity between the reflectionless and echo mode spectra. For asymptotic modes satisfying $|\mathrm{Re}\omega|\gg|\mathrm{Im}\omega|$, it is shown that these two spectra exhibit a strong resemblance, featuring an approximately uniform distribution parallel to the real frequency axis with the same spacing between successive modes. In particular, it was shown that the real parts of asymptotic QNMs coincide with those of RSMs. While echo modes typically possess non-vanishing imaginary parts, the RSMs of Damour-Solodukhin wormholes mainly lie close to the real frequency axis, where the distance between the RSMs and the real frequency axis measures the degree of deviation from a perfectly symmetric Damour-Solodukhin. Compared with echo modes, this property of RSMs leads to more substantial amplitudes in the time-domain waveforms, and consequently stronger echoes. This is demonstrated by numerically evaluating the waveforms associated with the two types of modes for a given identical source. We carry out the derivations using two complementary approaches, based on the scattering matrix and the Green’s function.
@@ Acknowledgements
Appendix A Flux conservation and the nonvanishing imaginary parts of the reflectionless modes